OpenStudy (anonymous):

Given that P(A) = 0.5, P(B) = 0.6, and P(A and B) = 0.30, determine P(A|B)

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

so write \[\frac{.3}{.6}=\frac{1}{2}\]